OpenStudy (anonymous):

Find a polynomial function that has the given zeros. (There are many correct answers.) 0, 4, 10

OpenStudy (anonymous):

x(x-4)(x-10)=0 expand this

OpenStudy (anonymous):

\[(x^2-4x)(x-10)=0 ......(x^3-10x^2-4x^2+40x)=0\]

OpenStudy (anonymous):

now I see what I did wrong, thank you so much.

OpenStudy (anonymous):

:) YW